OpenStudy (anonymous):

Tony receives $10 a week for his allowance. At the beginning of the month he had $30 in a bank account. He spent $7.50 to buy a movie ticket, $3.25 for popcorn, and $60 on a new video game. After 4 weeks, how much money did Tony have?

OpenStudy (radar):

Forget the spending for now lets figure the total money that Tony could have if he didn't spend. $30+4($10) = 30 + 40 =$70. But Tony spent money, he spent $7.50+$3.25+$60.00=$70.75 The question is how much did Tony. Bad news, Tony spent more than he had: $70.00-$70.75= $-.75

OpenStudy (anonymous):

Thank you very much. You are very smart

OpenStudy (radar):

Keep at it, practice will end up making it easier. No problem